# Initial value space of the fourth order Painlevé system with $A_{5}^{(1)}\times A_{1}^{(1)}$ symmetry

### v2

@@ 3 Accessible singularities
$C_{2}^{(01)}\cap\left(C_{3}^{(01)}\cup C_{3}^{(02)}\right)=\emptyset,\quad C_{2}^{(01)}\cap\left(C_{3}^{(11)}\cup C_{3}^{(12)}\right)=\emptyset,\quad C_{2}^{(01)}\cap\left(C_{3}^{(21)}\cup C_{3}^{(22)}\right)=\emptyset,$
$C_{2}^{(12)}\cap\left(C_{3}^{(01)}\cup C_{3}^{(02)}\right)=\emptyset,\quad C_{2}^{(12)}\cap\left(C_{3}^{(11)}\cup C_{3}^{(12)}\right)=\emptyset,\quad C_{2}^{(12)}\cap\left(C_{3}^{(21)}\cup C_{3}^{(22)}\right)=\emptyset,$
$C_{2}^{(22)}\cap\left(C_{3}^{(01)}\cup C_{3}^{(02)}\right)=\emptyset,\quad C_{2}^{(22)}\cap\left(C_{3}^{(11)}\cup C_{3}^{(12)}\right)=\emptyset,\quad C_{2}^{(22)}\cap\left(C_{3}^{(21)}\cup C_{3}^{(22)}\right)=\emptyset.$
Note that
Moreover, the accessible singularity $C_{2}^{(01)}$ is transformed to
$\left\{\left(\xi,\zeta^{(1)},t\right)\bigm{|}\xi_{0}-\xi_{1}=\zeta_{0}^{(1)}=\zeta_{2}^{(1)}=0,\ \xi_{1}\neq 0,\ \zeta_{1}^{(1)}\neq 0\right\}=\left\{\left([\xi_{1}:\xi_{1}:\xi_{2}],[0:1:0],t\right)\bigm{|}\xi_{1}\neq 0\right\}$
via relation (2.1), from which we obtain $C_{2}^{(01)}\cap C_{2}^{(12)}=\emptyset$. We also obtain
$C_{2}^{(01)}\cap C_{2}^{(22)}=\emptyset,\quad C_{2}^{(12)}\cap C_{2}^{(22)}=\emptyset$
in a similar manner.
We next consider the rest three accessible singularities. Using the homogeneous coordinates (2.2), we have
$C_{1}^{(01)}$ $=\left\{\left(\xi,\zeta^{(0)},t\right)\bigm{|}\xi_{1}-\xi_{2}=\zeta_{0}^{(0)}=\zeta_{1}^{(0)}+\zeta_{2}^{(0)}=0,\ \xi_{0}\neq 0,\ \zeta_{1}^{(0)}\neq 0\right\}$
$=\left\{\left([\xi_{0}:\xi_{1}:\xi_{1}],[0:1:-1],t\right)\bigm{|}\xi_{0}\neq 0\right\}.$
Under the condition $\xi_{1}\neq 0$, via relation (2.1), the accessible singularity $C_{1}^{(01)}$ is transformed to
$\left\{\left(\xi,\zeta^{(1)},t\right)\bigm{|}\xi_{1}-\xi_{2}=\zeta_{0}^{(1)}=\zeta_{1}^{(1)}=0,\ \xi_{0}\neq 0,\ \xi_{1}\neq 0,\ \zeta_{2}^{(1)}\neq 0\right\}$
$=\left\{\left([\xi_{0}:\xi_{1}:\xi_{1}],[0:0:1],t\right)\bigm{|}\xi_{0}\neq 0,\ \xi_{1}\neq 0\right\},$
which is contained in $C_{2}^{(12)}$. Under the condition $\xi_{1}=0$, $C_{1}^{(01)}$ is equivalent to
$\left\{\left(\xi,\zeta^{(0)},t\right)\bigm{|}\xi_{1}=\xi_{2}=\zeta_{0}^{(0)}=\zeta_{1}^{(0)}+\zeta_{2}^{(0)}=0,\ \xi_{0}\neq 0,\ \zeta_{1}^{(0)}\neq 0\right\}=\left\{\left([1:0:0],[0:1:-1],t\right)\right\},$
which is contained in $C_{3}^{(01)}\cup C_{3}^{(02)}$. Hence we obtain
$C_{1}^{(01)}\subset C_{2}^{(12)}\sqcup\left(C_{3}^{(01)}\cup C_{3}^{(02)}\right).$
We also obtain
$C_{1}^{(12)}\subset C_{2}^{(22)}\sqcup\left(C_{3}^{(11)}\cup C_{3}^{(12)}\right),\quad C_{1}^{(22)}\subset C_{2}^{(01)}\sqcup\left(C_{3}^{(21)}\cup C_{3}^{(22)}\right)$
in a similar manner.
In the last, we consider the accessible singularities $P_{j}^{(kl)}$. Using the homogeneous coordinates (2.2), we have
$=\left\{\left([1:0:t],[0:1:0],t\right)\right\}.$
Recall that $P_{3}^{(01)}$ is equivalent to
$=\left\{\left([1:0:t],[0:1:0],t\right)\right\}.$
We can see that $P_{3}^{(01)}=P_{3}^{(21)}$ is not contained in $C$. We can prove for the other accessible singularities in a similar manner. ∎
Lemma 3.2, Lemma 3.3 and Lemma 3.4 imply the following proposition.
Proposition 3.5.
All of the accessible singularities of the system $\mathcal{H}$ is reduced to
$C_{2}^{(01)},\ C_{3}^{(01)},\ C_{3}^{(02)},\ C_{3}^{(11)},\ C_{2}^{(12)},\ C_{3}^{(12)},\ C_{3}^{(21)},\ C_{2}^{(22)},\ C_{3}^{(22)}$ (3.3)
@@
$P_{3}^{(01)},\ P_{4}^{(01)},\ P_{6}^{(01)},\ P_{7}^{(01)},\ P_{3}^{(12)},\ P_{4}^{(12)},\ P_{6}^{(12)},\ P_{7}^{(12)},\ P_{3}^{(22)},\ P_{4}^{(22)},\ P_{6}^{(22)},\ P_{7}^{(22)}.$ (3.4)
4 Resolving the accessible singularities $C_{j}^{(kl)}$
In the previous section, we have determined the accessible singularities of the system $\mathcal{H}$. As will be seen below, there are infinitely many solutions passing through each accessible singularity. In this section, we resolve the accessible singularities (3.3) by introducing new coordinates which parametrize these solutions. We give the result of this section in advance.
Proposition 4.1.
Coordinates resolving the accessible singularities (3.3) are given as follows.
$C_{2}^{(01)}:\left(x_{2}^{[01]},y_{2}^{[01]},z_{2}^{[01]},w_{2}^{[01]}\right)=\left(-\{(q_{1}-1)p_{1}-\alpha_{0}\}p_{1},q_{2},\dfrac{1}{p_{1}},p_{2}\right).$
$C_{3}^{(01)}:\left(x_{3}^{[01]},y_{3}^{[01]},z_{3}^{[01]},w_{3}^{[01]}\right)=\left(-(q_{1}p_{1}+q_{2}p_{2}-\alpha_{5}+\eta)p_{1},q_{2}p_{1},\dfrac{1}{p_{1}},\dfrac{p_{2}}{p_{1}}\right).$
$C_{3}^{(02)}:\left(x_{3}^{[02]},y_{3}^{[02]},z_{3}^{[02]},w_{3}^{[02]}\right)=\left(q_{1}p_{2},-(q_{1}p_{1}+q_{2}p_{2}-\alpha_{5}+\eta)p_{2},\dfrac{p_{1}}{p_{2}},\dfrac{1}{p_{2}}\right).$
$C_{3}^{(11)}:\left(x_{3}^{[11]},y_{3}^{[11]},z_{3}^{[11]},w_{3}^{[11]}\right)=\left(-\left(q_{1}^{(1)}p_{1}^{(1)}+q_{2}^{(1)}p_{2}^{(1)}-\alpha_{1}+\eta\right)p_{1}^{(1)},q_{2}^{(1)}p_{1}^{(1)},\dfrac{1}{p_{1}^{(1)}},\dfrac{p_{2}^{(1)}}{p_{1}^{(1)}}\right).$
$C_{2}^{(12)}:\left(x_{2}^{[12]},y_{2}^{[12]},z_{2}^{[12]},w_{2}^{[12]}\right)=\left(q_{1}^{(1)},-\left\{\left(q_{2}^{(1)}-1\right)p_{2}^{(1)}-\alpha_{2}\right\}p_{2}^{(1)},p_{1}^{(1)},\dfrac{1}{p_{2}^{(1)}}\right).$
$C_{3}^{(12)}:\left(x_{3}^{[12]},y_{3}^{[12]},z_{3}^{[12]},w_{3}^{[12]}\right)=\left(q_{1}^{(1)}p_{2}^{(1)},-\left(q_{1}^{(1)}p_{1}^{(1)}+q_{2}^{(1)}p_{2}^{(1)}-\alpha_{1}+\eta\right)p_{2}^{(1)},\dfrac{p_{1}^{(1)}}{p_{2}^{(1)}},\dfrac{1}{p_{2}^{(1)}}\right).$
$C_{3}^{(21)}:\left(x_{3}^{[21]},y_{3}^{[21]},z_{3}^{[21]},w_{3}^{[21]}\right)=\left(-\left(q_{1}^{(2)}p_{1}^{(2)}+q_{2}^{(2)}p_{2}^{(2)}-\alpha_{3}+\eta\right)p_{1}^{(2)},q_{2}^{(2)}p_{1}^{(2)},\dfrac{1}{p_{1}^{(2)}},\dfrac{p_{2}^{(2)}}{p_{1}^{(2)}}\right).$
$C_{2}^{(22)}:\left(x_{2}^{[22]},y_{2}^{[22]},z_{2}^{[22]},w_{2}^{[22]}\right)=\left(q_{1}^{(2)},-\left\{\left(q_{2}^{(2)}-\frac{1}{t}\right)p_{2}^{(2)}-\alpha_{4}\right\}p_{2}^{(2)},p_{1}^{(2)},\frac{1}{p_{2}^{(2)}}\right).$
$C_{3}^{(22)}:\left(x_{3}^{[22]},y_{3}^{[22]},z_{3}^{[22]},w_{3}^{[22]}\right)=\left(q_{1}^{(2)}p_{2}^{(2)},-\left(q_{1}^{(2)}p_{1}^{(2)}+q_{2}^{(2)}p_{2}^{(2)}-\alpha_{3}+\eta\right)p_{2}^{(2)},\dfrac{p_{1}^{(2)}}{p_{2}^{(2)}},\dfrac{1}{p_{2}^{(2)}}\right).$
Namely, the system on each coordinate $\left(x_{j}^{[kl]},y_{j}^{[kl]},z_{j}^{[kl]},w_{j}^{[kl]}\right)$ is described as a polynomial Hamiltonian system.
In this proposition, we denote the coordinate resolving $C_{j}^{(kl)}$ by $\left(x_{j}^{[kl]},y_{j}^{[kl]},z_{j}^{[kl]},w_{j}^{[kl]}\right)$. This proposition is shown in two ways. One is the blowing up method and another is that in terms of formal power series. Note that the latter method will be used in the next section to resolve $P_{j}^{(kl)}$. Because we can prove in the same way for all of the accessible singularities, we consider only $C_{2}^{(01)}=\left\{q_{1}-1=p_{1}^{(01)}=p_{2}^{(01)}=0\right\}$ here. The system $\mathcal{H}$ on $W_{01}\times B$ is described as
$\begin{split}t(t-1)\frac{dq_{1}}{dt}&=(\alpha_{1}+\eta)q_{1}^{2}+\alpha_{3}q_{1}q_{2}+\{(\alpha_{0}+\alpha_{5}-\eta)t-(\alpha_{0}+\alpha_{1}+\eta)\}q_{1}-\alpha_{3}q_{2}-(\alpha_{5}-\eta)t\\ &\quad+\frac{(q_{1}-1)(q_{1}+q_{2})(q_{2}-t)p_{2}^{(01)}+2q_{1}(q_{1}-1)(q_{1}-t)}{p_{1}^{(01)}},\\ t(t-1)\frac{dq_{2}}{dt}&=\alpha_{1}q_{1}q_{2}+(\alpha_{3}+\eta)q_{2}^{2}-\alpha_{1}tq_{1}-\{(\alpha_{3}+\alpha_{4}+\eta-1)t-(\alpha_{4}+\alpha_{5}-\eta-1)\}q_{2}-(\alpha_{5}-\eta)t\\ &\quad+\frac{2q_{2}(q_{2}-1)(q_{2}-t)p_{2}^{(01)}+(q_{1}-1)(q_{1}+q_{2})(q_{2}-t)}{p_{1}^{(01)}},\\ t(t-1)\frac{dp_{1}^{(01)}}{dt}&=\alpha_{1}\eta\left(p_{1}^{(01)}\right)^{2}+\{2(\alpha_{1}+\eta)q_{1}+(\alpha_{1}+\alpha_{3})q_{2}+(\alpha_{0}-\alpha_{1}+\alpha_{5}-\eta)t-(\alpha_{0}+\alpha_{1}+\eta)\}p_{1}^{(01)}\\ &\quad+(2q_{1}+q_{2}-1)(q_{2}-t)p_{2}^{(01)}+3q_{1}^{2}-2(t+1)q_{1}+t,\\ t(t-1)\frac{dp_{2}^{(01)}}{dt}&=\eta\left(\alpha_{1}p_{2}^{(01)}-\alpha_{3}\right)p_{1}^{(01)}+\alpha_{1}(q_{2}-t)\left(p_{2}^{(01)}\right)^{2}\\ &\quad+\{(\alpha_{1}+2\eta)q_{1}-(\alpha_{3}+2\eta)q_{2}-\alpha_{1}t+\alpha_{2}+\alpha_{3}\}p_{2}^{(01)}-\alpha_{3}(q_{1}-1)\\ &\quad+\frac{\left\{2q_{1}q_{2}-2q_{2}^{2}-2tq_{1}+(t+1)q_{2}\right\}\left(p_{2}^{(01)}\right)^{2}+\left\{2q_{1}q_{2}+2q_{2}^{2}-(t+1)q_{1}+2q_{2}\right\}p_{2}^{(01)}}{p_{1}^{(01)}}.\end{split}$ (4.1)
4.1 Blowing up
Let $\left(u_{1},u_{2},v_{1},v_{2}\right)$, $\left(u^{\prime}_{1},u^{\prime}_{2},v^{\prime}_{1},v^{\prime}_{2}\right)$ and $\left(u^{\prime\prime}_{1},u^{\prime\prime}_{2},v^{\prime\prime}_{1},v^{\prime\prime}_{2}\right)$ be affine coordinates of $\mathbb{C}^{4}$ defined by
$\begin{split}&q_{1}-1=u_{1}v_{1},\quad q_{2}=u_{2},\quad p_{1}^{(01)}=v_{1},\quad p_{2}^{(01)}=v_{1}v_{2},\\ &q_{1}-1=u^{\prime}_{1},\quad q_{2}=u^{\prime}_{2},\quad p_{1}^{(01)}=u^{\prime}_{1}v^{\prime}_{1},\quad p_{2}^{(01)}=u^{\prime}_{1}v^{\prime}_{2},\\ &q_{1}-1=u^{\prime\prime}_{1}v^{\prime\prime}_{2},\quad q_{2}=u^{\prime\prime}_{2},\quad p_{1}^{(01)}=v^{\prime\prime}_{1}v^{\prime\prime}_{2},\quad p_{2}^{(01)}=v^{\prime\prime}_{2}.\end{split}$ (4.2)
On these coordinates, the accessible singularity $C_{2}^{(01)}$ corresponds to
$E_{1}=\left\{v_{1}=0\right\},\quad E^{\prime}_{1}=\left\{u^{\prime}_{1}=0\right\},\quad E^{\prime\prime}_{1}=\left\{v^{\prime\prime}_{2}=0\right\}.$
The union $E_{1}\cup E^{\prime}_{1}\cup E^{\prime\prime}_{1}$ is called the exceptional divisor and $E_{1}\cup E^{\prime}_{1}\cup E^{\prime\prime}_{1}\simeq\mathbb{C}^{1}\times\mathbb{P}^{2}$. System (4.1) on the coordinate $(u_{1},u_{2},v_{1},v_{2})$ is described as
$\frac{du_{1}}{dt}=\frac{\alpha_{0}-u_{1}}{tv_{1}}+U_{1},\quad\frac{du_{2}}{dt}=U_{2},\quad\frac{dv_{1}}{dt}=V_{1},\quad\frac{dv_{2}}{dt}=V_{2},$ (4.3)
where $U_{1},U_{2},V_{1},V_{2}\in\mathbb{C}(t)[u_{1},u_{2},v_{1},v_{2}]$ and its accessible singularity is given by
$D_{1}=\left\{\alpha_{0}-u_{1}=v_{1}=0\right\}\subset E_{1}.$
System (4.1) on the coordinate $(u^{\prime}_{1},u^{\prime}_{2},v^{\prime}_{1},v^{\prime}_{2})$ is described as
$\frac{du^{\prime}_{1}}{dt}=\dfrac{U^{\prime}_{1}}{v^{\prime}_{1}},\quad\frac{du^{\prime}_{2}}{dt}=\dfrac{U^{\prime}_{2}}{v^{\prime}_{1}},\quad\frac{dv^{\prime}_{1}}{dt}=\dfrac{1-\alpha_{0}v^{\prime}_{1}}{tu^{\prime}_{1}}+V^{\prime}_{1},\quad\frac{dv^{\prime}_{2}}{dt}=\dfrac{(t-1)\left(1-\alpha_{0}v^{\prime}_{1}\right)v^{\prime}_{2}+u^{\prime}_{1}V^{\prime}_{2}}{t(t-1)u^{\prime}_{1}v^{\prime}_{1}},$
@@ 4.1 Blowing up
where $U^{\prime}_{1},U^{\prime}_{2},V^{\prime}_{1},V^{\prime}_{2}\in\mathbb{C}(t)\left[u^{\prime}_{1},u^{\prime}_{2},v^{\prime}_{1},v^{\prime}_{2}\right]$ and its accessible singularity is given by
$D^{\prime}_{1}=\left\{1-\alpha_{0}v^{\prime}_{1}=u^{\prime}_{1}=0\right\}\subset E^{\prime}_{1}.$
System (4.1) on the coordinate $(u^{\prime\prime}_{1},u^{\prime\prime}_{2},v^{\prime\prime}_{1},v^{\prime\prime}_{2})$ is described as
$\frac{du^{\prime\prime}_{1}}{dt}=\dfrac{\left(t-1\right)\left(\alpha_{0}v^{\prime\prime}_{1}-u^{\prime\prime}_{1}\right)+v^{\prime\prime}_{2}U^{\prime\prime}_{1}}{t\left(t-1\right)v^{\prime\prime}_{1}v^{\prime\prime}_{2}},\quad\frac{du^{\prime\prime}_{2}}{dt}=\dfrac{U^{\prime\prime}_{2}}{v^{\prime\prime}_{1}},\quad\frac{dv^{\prime\prime}_{1}}{dt}=V^{\prime\prime}_{1},\quad\frac{dv^{\prime\prime}_{2}}{dt}=\dfrac{1-t+v^{\prime\prime}_{2}V^{\prime\prime}_{2}}{t\left(t-1\right)v^{\prime\prime}_{1}},$
where $U^{\prime\prime}_{1},U^{\prime\prime}_{2},V^{\prime\prime}_{1},V^{\prime\prime}_{2}\in\mathbb{C}(t)\left[u^{\prime\prime}_{1},u^{\prime\prime}_{2},v^{\prime\prime}_{1},v^{\prime\prime}_{2}\right]$ and its accessible singularity is given by
$D^{\prime\prime}_{1}=\left\{\alpha_{0}v^{\prime\prime}_{1}-u^{\prime\prime}_{1}=v^{\prime\prime}_{2}=0,\ v^{\prime\prime}_{1}\neq 0\right\}\subset E^{\prime\prime}_{1}.$
Relation (4.2) implies
$\left(u^{\prime}_{1},u^{\prime}_{2},v^{\prime}_{1},v^{\prime}_{2}\right)=\left(u_{1}v_{1},u_{2},\frac{1}{u_{1}},\frac{v_{2}}{u_{1}}\right),\quad\left(u^{\prime\prime}_{1},u^{\prime\prime}_{2},v^{\prime\prime}_{1},v^{\prime\prime}_{2}\right)=\left(\frac{u_{1}}{v_{2}},u_{2},\frac{1}{v_{2}},v_{1}v_{2}\right),$
from which we have $D^{\prime}_{1}=D_{1}$ and $D^{\prime\prime}_{1}\subset D_{1}$. Note that $D_{1}\simeq\mathbb{C}^{2}$. Hence it is enough to consider system (4.3) and its accessible singularity $D_{1}$.
Let $\left(x_{2}^{[01]},y_{2}^{[01]},z_{2}^{[01]},w_{2}^{[01]}\right)$ and $(x^{\prime}_{2},y^{\prime}_{2},z^{\prime}_{2},w^{\prime}_{2})$ be affine coordinates of $\mathbb{C}^{4}$ defined by
$\begin{split}&\alpha_{0}-u_{1}=x_{2}^{[01]}z_{2}^{[01]},\quad u_{2}=y_{2}^{[01]},\quad v_{1}=z_{2}^{[01]},\quad v_{2}=w_{2}^{[01]},\\ &\alpha_{0}-u_{1}=x^{\prime}_{2},\quad u_{2}=y^{\prime}_{2},\quad v_{1}=x^{\prime}_{2}z^{\prime}_{2},\quad v_{2}=w^{\prime}_{2}.\end{split}$ (4.4)
On these coordinates, the accessible singularity $D_{1}$ corresponds to
$E_{2}=\left\{z_{2}^{[01]}=0\right\},\quad E^{\prime}_{2}=\left\{x^{\prime}_{2}=0\right\}.$
Note that $E_{2}\cup E^{\prime}_{2}\simeq\mathbb{C}^{2}\times\mathbb{P}^{1}$. System (4.3) on the coordinate $\left(x_{2}^{[01]},y_{2}^{[01]},z_{2}^{[01]},w_{2}^{[01]}\right)$ is described as
$\frac{dx_{2}^{[01]}}{dt}=X_{2},\quad\frac{dy_{2}^{[01]}}{dt}=Y_{2},\quad\frac{dz_{2}^{[01]}}{dt}=Z_{2},\quad\frac{dw_{2}^{[01]}}{dt}=W_{2},$ (4.5)
where $X_{2},Y_{2},Z_{2},W_{2}\in\mathbb{C}(t)\left[x_{2}^{[01]},y_{2}^{[01]},z_{2}^{[01]},w_{2}^{[01]}\right]$ and it does not have any singularity. System (4.3) on the coordinate $\left(x^{\prime}_{2},y^{\prime}_{2},z^{\prime}_{2},w^{\prime}_{2}\right)$ is described as
$\frac{dx^{\prime}_{2}}{dt}=\dfrac{1-t+z^{\prime}_{2}X^{\prime}_{2}}{t(t-1)z^{\prime}_{2}},\quad\frac{dy^{\prime}_{2}}{dt}=Y^{\prime}_{2},\quad\frac{dz^{\prime}_{2}}{dt}=Z^{\prime}_{2},\quad\frac{dw^{\prime}_{2}}{dt}=W^{\prime}_{2},$
where $X^{\prime}_{2},Y^{\prime}_{2},Z^{\prime}_{2},W^{\prime}_{2}\in\mathbb{C}(t)\left[x^{\prime}_{2},y^{\prime}_{2},z^{\prime}_{2},w^{\prime}_{2}\right]$ and it does not have any accessible singularity. Note that the condition $z^{\prime}_{2}=0$ corresponds to $x^{[01]}_{2}$$\infty$ and the solution satisfying $z^{\prime}_{2}=0$ does not pass through $\Sigma_{\eta}^{0}$. Hence it is enough to consider system (4.5).

### v3

@@ 5.1 Blowing up
$\left\{u_{12}+u_{22}v_{22}-\alpha_{5}+\eta=v_{12}=0\right\}$ (5.4)
respectively. System (3.6) on the coordinate $(u_{12}^{\prime},u_{22}^{\prime},v_{12}^{\prime},v_{22}^{\prime})$ is described as
$\begin{split}&t(t-1)v_{12}^{\prime}du_{12}^{\prime}-U_{12}^{\prime}dt=0,\\ &t(t-1)u_{12}^{\prime}v_{12}^{\prime}du_{22}^{\prime}-U_{22}^{\prime}dt=0,\\ &t(t-1)u_{12}^{\prime}dv_{12}^{\prime}-V_{12}^{\prime}dt=0,\\ &t(t-1)v_{12}^{\prime}dv_{22}^{\prime}-V_{22}^{\prime}dt=0,\end{split}$ (5.5)
where $U_{12}^{\prime},U_{22}^{\prime},V_{12}^{\prime},V_{22}^{\prime}\in\mathbb{C}\left[u_{12}^{\prime},u_{22}^{\prime},v_{12}^{\prime},v_{22}^{\prime},t\right]$. The singularity and the accessible singularity of system (5.5) on $D_{1}^{1}$ are given by
$\left\{u_{12}^{\prime}=0\right\}\cup\left\{v_{12}^{\prime}=0\right\}$
@@
$\left\{u^{\prime}_{12}=u_{22}^{\prime}v_{22}^{\prime}+(-\alpha_{5}+\eta)v_{12}^{\prime}+1=0,\ v_{12}^{\prime}\neq 0\right\}\cup\left\{u^{\prime}_{12}=u_{22}^{\prime}-1=v^{\prime}_{12}=v_{22}^{\prime}+1=0\right\}$ (5.6)
respectively. System (3.6) on the coordinate $(u_{12}^{\prime\prime},u_{22}^{\prime\prime},v_{12}^{\prime\prime},v_{22}^{\prime\prime})$ is described as
$t(t-1)u_{22}^{\prime\prime}v_{12}^{\prime\prime}du_{12}^{\prime\prime}-U_{12}^{\prime\prime}dt=0,$
$t(t-1)v_{12}^{\prime\prime}du_{22}^{\prime\prime}-U_{22}^{\prime\prime}dt=0,$
$t(t-1)u_{22}^{\prime\prime}dv_{12}^{\prime\prime}-V_{12}^{\prime\prime}dt=0,$
$t(t-1)v_{12}^{\prime\prime}dv_{22}^{\prime\prime}-V_{22}^{\prime\prime}dt=0,$
@@
where $U_{12}^{\prime\prime},U_{22}^{\prime\prime},V_{12}^{\prime\prime},V_{22}^{\prime\prime}\in\mathbb{C}\left[u^{\prime\prime}_{12},u^{\prime\prime}_{22},v^{\prime\prime}_{12},v^{\prime\prime}_{22},t\right]$. The singularity and the accessible singularity of this system on $D_{1}^{1}$ are given by
$\left\{u_{22}^{\prime\prime}=0\right\}\cup\left\{v_{12}^{\prime\prime}=0\right\}$
and
$\left\{u^{\prime\prime}_{12}+(-\alpha_{5}+\eta)v_{12}^{\prime\prime}+v_{22}^{\prime\prime}=u_{22}^{\prime\prime}=0,\ v^{\prime\prime}_{12}\neq 0\right\}\cup\left\{u^{\prime\prime}_{12}-1=u^{\prime\prime}_{22}=v_{12}^{\prime\prime}=v_{22}^{\prime\prime}+1=0\right\}$ (5.7)
respectively. Relation (5.1) implies
$\left(u^{\prime}_{12},u^{\prime}_{22},v^{\prime}_{12},v^{\prime}_{22}\right)=\left(u_{12}v_{12},\frac{u_{22}}{u_{12}},\frac{1}{u_{12}},v_{22}\right),$
$\left(u^{\prime\prime}_{12},u^{\prime\prime}_{22},v^{\prime\prime}_{12},v^{\prime\prime}_{22}\right)=\left(\frac{u_{12}}{u_{22}},u_{22}v_{12},\frac{1}{u_{22}},v_{22}\right),$
$\left(u^{\prime}_{12},u^{\prime}_{22},v^{\prime}_{12},v^{\prime}_{22}\right)=\left(u_{12}^{\prime\prime}u_{22}^{\prime\prime},\frac{1}{u_{12}^{\prime\prime}},\frac{v_{12}^{\prime\prime}}{u_{11}^{\prime\prime}},v_{22}^{\prime\prime}\right),$
from which the first components (5.6) and (5.7) are both contained in (5.4) and the second components (5.6) and (5.7) are equal. We can consider the other accessible singularities $C_{1}^{(01)}$, $C_{3}^{(02)}$ and $C_{2}^{(12)}$ in a similar manner. Consequently, we obtain new accessible singularities
$\begin{split}C_{1}^{1}&=\left\{u_{11}-\alpha_{2}=v_{11}=0\right\}\cup\left\{u_{12}+u_{22}v_{22}-\alpha_{5}+\eta=v_{12}=0\right\}\\ &\quad\cup\left\{u_{13}v_{13}+u_{23}-\alpha_{5}+\eta=v_{23}=0\right\}\cup\left\{u_{24}-\alpha_{2}=v_{24}=0\right\},\\ P_{1}^{1}&=\left\{u^{\prime}_{12}=u_{22}^{\prime}-1=v^{\prime}_{12}=v_{22}^{\prime}+1=0\right\}\end{split}$ (5.8)
on $D_{1}^{1}$. We extend the system $\mathcal{H}$ to that on $\left(\Sigma_{\eta}\cup D_{1}^{1}\right)\times B$ and denote it by $\mathcal{H}_{1}^{1}$ temporarily. Then there exist transversal leaves corresponding to the solutions to $\mathcal{H}_{1}^{1}$ passing through $C_{1}^{1}\cup P_{1}^{1}$. Moreover, all of the leaves corresponding to the solutions to $\mathcal{H}_{1}^{1}$ passing through the singularity $D_{1}^{1}\setminus\left(C_{1}^{1}\cup P_{1}^{1}\right)$ are vertical leaves. This fact can be shown in a similar manner as in $D^{0}\setminus(C^{0}\cup P^{0})$. We do not state its detail here.
We next consider the blowing up at the accessible singularity $C_{1}^{1}$. We set eight affine coordinates of $\mathbb{C}^{4}$ as
$\alpha_{2}-u_{11}=x_{1}^{[01]}z_{1}^{[01]},\quad u_{21}=y_{1}^{[01]},\quad v_{11}=z_{1}^{[01]},\quad v_{21}=w_{1}^{[01]},$
$\alpha_{2}-u_{11}=x^{\prime}_{1},\quad u_{21}=y^{\prime}_{1},\quad v_{11}=x^{\prime}_{1}z^{\prime}_{1},\quad v_{21}=w^{\prime}_{1},$
$-(u_{12}+u_{22}v_{22}-\alpha_{5}+\eta)=x_{3}^{[01]}z_{3}^{[01]},\quad u_{22}=y_{3}^{[01]},\quad v_{12}=z_{3}^{[01]},\quad v_{22}=w_{3}^{[01]},$
$-(u_{12}+u_{22}v_{22}-\alpha_{5}+\eta)=x_{31}^{\prime},\quad u_{22}=y_{31}^{\prime},\quad v_{12}=x_{31}^{\prime}z_{31}^{\prime},\quad v_{22}=w_{31}^{\prime},$
$u_{13}=x_{3}^{[02]},\quad-(u_{13}v_{13}+u_{23}-\alpha_{5}+\eta)=y_{3}^{[02]}w_{3}^{[02]},\quad v_{13}=z_{3}^{[02]},\quad v_{23}=w_{3}^{[02]},$
$u_{13}=x_{32}^{\prime},\quad-(u_{13}v_{13}+u_{23}-\alpha_{5}+\eta)=y_{32}^{\prime},\quad v_{13}=z_{32}^{\prime},\quad v_{23}=y_{32}^{\prime}w_{32}^{\prime},$
$u_{14}=x_{2}^{[12]},\quad\alpha_{2}-u_{24}=y_{2}^{[12]}w_{2}^{[12]},\quad v_{14}=z_{2}^{[12]},\quad v_{24}=w_{2}^{[12]},$
$u_{14}=x_{2}^{\prime},\quad\alpha_{2}-u_{24}=y_{2}^{\prime},\quad v_{14}=z_{2}^{\prime},\quad v_{24}=y_{2}^{\prime}w_{2}^{\prime}.$
On these coordinates, the accessible singularity $C_{1}^{1}$ corresponds to
$\begin{split}D_{1}^{2}&=\left\{z_{1}^{[01]}=0\right\}\cup\left\{x^{\prime}_{1}=0\right\}\cup\left\{z_{3}^{[01]}=0\right\}\cup\left\{x^{\prime}_{31}=0\right\}\\ &\quad\cup\left\{w_{3}^{[02]}=0\right\}\cup\left\{y^{\prime}_{32}=0\right\}\cup\left\{w_{2}^{[12]}=0\right\}\cup\left\{y^{\prime}_{2}=0\right\}.\end{split}$ (5.9)
The union $D_{1}^{2}$ is connected and called an exceptional divisor. We consider the accessible singularity $\left\{u_{12}+u_{22}v_{22}-\alpha_{5}+\eta=v_{12}=0\right\}$. System (5.3) on the coordinate $\left(x_{3}^{[01]},y_{3}^{[01]},z_{3}^{[01]},w_{3}^{[01]}\right)$ is described as
$\begin{split}&t(t-1)dx_{3}^{[01]}-X_{3}dt=0,\\ &t(t-1)dy_{3}^{[01]}-Y_{3}dt=0,\\ &t(t-1)dz_{3}^{[01]}-Z_{3}dt=0,\\ &t(t-1)dw_{3}^{[01]}-W_{3}dt=0,\end{split}$ (5.10)